OpenStudy (anonymous):

PLEASE HELP!!! Solving Equations with Variables on Both Sides Solve for x: −2(x + 3) = −2(x + 1) − 4

OpenStudy (anonymous):

expand both sides, −2(x + 3)=-2x-6 −2(x + 1) − 4=-2x-2-4=-2x-6 then you get this -2x-6=-2x-6 which tells you x=x. This means there are infinite solutions to this problem. So x can be any number.

OpenStudy (anonymous):

so no solution ?

OpenStudy (anonymous):

2 3 All real numbers No solution

OpenStudy (anonymous):

thoses are the options ^^

OpenStudy (anonymous):

All real numbers. Value of x can be anything since x=x

OpenStudy (anonymous):

ohhh ok thanks